# Differentially Private Multi-objective Selection: Pareto and Aggregation Approaches

### v1

@@ Abstract.
Differentially private selection mechanisms are fundamental building blocks for privacy-preserving data analysis. While numerous mechanisms exist for single-objective selection, many real-world applications require optimizing multiple competing objectives simultaneously. We present two novel mechanisms for differentially private multi-objective selection: PrivPareto and PrivAgg. PrivPareto uses a novel Pareto score to identify solutions near the Pareto frontier, while PrivAgg enables privacy-preserving weighted aggregation of multiple objectives. Both mechanisms support global and local sensitivity approaches, with comprehensive theoretical analysis showing how to compose sensitivities of multiple utility functions. We demonstrate the practical applicability through two real-world applications: cost-sensitive decision tree construction and multi-objective influential node selection in social networks. The experimental results showed that our local sensitivity-based approaches achieve significantly better utility compared to global sensitivity approaches across both applications and both Pareto and Aggregation approaches. Moreover, the local sensitivity-based approaches are able to perform well with typical privacy budget values $\epsilon\in[0.01,1]$ in most experiments.
@@ 1. Introduction
Data analysis is a fundamental tool for decision making in many areas, such as healthcare, finance, and marketing. Data analysts can use data to extract valuable insights, make predictions, and optimize decisions and publish their findings. On the other hand, a malicious adversary can use the published data to infer sensitive information about individuals (Henriksen-Bulmer and Jeary, 2016; Hamza et al., 2013; Narayanan and Shmatikov, 2016).
@@
Differential privacy has been widely applied to the problem of single-objective private selection. In this problem, the goal is to select an element from a set of candidates that maximizes a single objective function. Formally, given a set of candidates $\mathcal{R}$ and a utility function $u:\mathcal{R}\times\mathcal{X}\rightarrow$ that takes a candidate $r\in\mathcal{R}$ and a dataset $x\in\mathcal{X}$ and outputs a score, the goal is to select an element $r\in\mathcal{R}$ that approximately maximizes $u(r)$ while satisfying differential privacy. The utility function $u$ is designed by the data analyst and measures how well a candidate $r$ satisfies the objective of a given task. For the rest of the paper, we will use utility function and objective function interchangeably.
Many algorithms have been proposed to solve this problem under differential privacy, the exponential mechanism (McSherry and Talwar, 2007), the report noisy max algorithm (Dwork et al., 2014), the permute-and-flip mechanism (McKenna and Sheldon, 2020) and the local dampening mechanism (de Farias et al., 2020; Farias et al., 2023). These mechanisms use the notion of sensitivity in two flavors, global sensitivity and local sensitivity, where these notions measure how much the utility function changes when a single individual is added or removed from a dataset.
Such algorithms are utterly important in the field of differential privacy since they are fundamental building blocks and can be composed to create more complex algorithms. The applications include private release of streams, histograms, databases and graphs (Zhu et al., 2017), data mining (Fletcher and Islam, 2019), machine learning (Ji et al., 2014; Liu et al., 2021) and social network analysis (Task and Clifton, 2012; Jiang et al., 2021).
@@
Thus, we define our problem statement of multi-objective private selection as follows. Given a set of candidates $\mathcal{R}$ and a set of utility functions $u_{1},u_{2},\ldots,u_{m}:\mathcal{R}\times\mathcal{X}\rightarrow$, the goal is to select an element $r\in\mathcal{R}$ that maximizes $\{u_{1}(c),u_{2}(c),\ldots,u_{m}(c)\}$ while satisfying differential privacy.
Note that the utility functions may be conflicting and there could be no candidate that maximizes all utility functions simultaneously. Instead, it is desired to find a set of candidates that are not dominated in all utility functions. This is are said to be Pareto optimal and the set of all Pareto optimal candidates is called the Pareto front. In this setting, we choose candidates that are the best in at least one of the objective functions. In the example of decision trees, the Pareto front would contains trees that maximize either the true positive rate or the true negative rate.
@@
To show the effectiveness and generality of our mechanisms, we conduct an extensive experimental evaluation on two distinct problems: (i) multi-objective influential node selection in social networks where we select the most influencial nodes in a graph based on multiple centrality measures, and (ii) cost-sensitive decision tree construction where we build decision trees that maximize the true positive rate and true negative rate simultaneously.
@@ 2.1. Basic Definitions
($\epsilon$-Differential Privacy (Dwork et al., 2006a; Dwork et al., 2006b)). A randomized algorithm $\mathcal{M}$ satisfies $\epsilon$-differential privacy if, for any two databases $x$ and $y$ such that $d(x,y)\leq 1$, and for any possible output $O$ of $\mathcal{M}$, the following inequality holds:
@@ 2.3. Sensitivity Functions
Computing local sensitivity $LS^{u}(x,t)$ or element local sensitivity $LS^{u}(x,t,r)$ can be infeasible is some cases where it is an NP-hard problem (Nissim et al., 2007; Zhang et al., 2015). To overcome this, we can build a computationally efficient function $\delta^{u}(x,t,r)$ that computes as an upper bound for $LS^{u}(x,t)$ or $LS^{u}(x,t,r)$, while staying below $\Delta u$. These functions are denoted as sensitivity functions. A sensitivity function needs to have the following signature: $\delta^{u}:\mathcal{X}\times\mathbb{N}\times\mathcal{R}\rightarrow$ and the function $\delta^{u}$ is said to be a sensitivity function for the utility function $u$. The $\delta^{u}(x,t,r)=\Delta u$, $\delta^{u}(x,t,r)=LS^{u}(x,t)$ and $\delta^{u}(x,t,r)=LS^{u}(x,t,r)$ are all sensitivity functions.
@@ 3. Related Work
The field of differentially private multi-objective selection remains unexplored. To the best of our knowledge, this work is the first to tackle this problem. Nonetheless, we reviewed the literature on related works in differential privacy for single-objective selection.
@@ 4. PrivPareto Mechanism
The choice of the single objective differentially private selection mechanism in the third step plays an important role on the accuracy of the our mechanism since each mechanism adds a different amount of noise to the output. These mechanism can use global or local sensitivity. Global sensitivity based algorithms are more frequent in the literature (McSherry and Talwar, 2007; McKenna and Sheldon, 2020). However, local sensitivity based mechanisms have been shown to perform better in applications where the local sensitivity is smaller than the global sensitivity (Farias et al., 2023) which is the case for the Pareto Score. Thus, we present the PrivPareto mechanism in two versions: a global sensitivity version and a local sensitivity version.
@@
(Global PrivPareto Mechanism) The global PrivPareto mechanism $PrivPareto_{global}(x,\epsilon,u_{1},\ldots,u_{m},\mathcal{R})$ takes as input a database $x\in\mathcal{X}$, the privacy budget $\epsilon$, the utility functions $u_{1},u_{2},\ldots,u_{m}:\mathcal{R}\times\mathcal{X}\rightarrow$, the set of candidate solutions $\mathcal{R}$, and outputs an approximate Pareto optimal solution $r\in\mathcal{R}$. The mechanism is defined as follows:
@@
The of the local sensitivity version of the PrivPareto mechanism is given in Definition 4.2:
@@
(Local PrivPareto Mechanism) The local PrivPareto mechanism $PrivPareto_{local}(x,\epsilon,u_{1},\ldots,u_{m},\delta^{u_{1}},\ldots,\delta^{u_{m}},\mathcal{R})$ takes as input a database $x\in\mathcal{X}$, the privacy budget $\epsilon$, the set of utility functions $u_{1},u_{2},\ldots,u_{m}:\mathcal{R}\times\mathcal{X}\rightarrow$, the admissible sensitivity functions $\delta^{u_{1}},\ldots,\delta^{u_{m}}$ of the utility functions, a set of candidate solutions $\mathcal{R}$, and outputs an approximate Pareto optimal solution $r\in\mathcal{R}$. The mechanism is defined as follows:
@@ 4.1. Pareto Score
(Pareto Score). Given a database $x\in\mathcal{X}$, a set of utility functions $u_{1},u_{2},\ldots,u_{m}:\mathcal{R}\times\mathcal{X}\rightarrow$, and a candidate solution $r\in\mathcal{R}$, the Pareto Score $PS(x,r)$ is defined as the number of candidates that dominate $r$ in the set $\mathcal{R}$:
@@
Consider a database $x$, a set of candidates =$\mathcal{R}=\{a,b,c,d,e\}$ and two utility functions $u_{1},u_{2}:\mathcal{R}\times\mathcal{X}\rightarrow$. The utility functions are defined as follows:
@@
Figure 1. Example Pareto scores for 5 elements and 2 utility functions. The left lower subspace of each candidate, delimited by the two dashed lines, contains the its dominated elements. For instance, candidate $a$ dominates $d$ and $e$.
@@ 4.2. Pareto Score Sensitivity
Consider a database $x$, a set of candidates =$\mathcal{R}=\{a,b,c\}$ and two utility functions $u_{1},u_{2}:\mathcal{R}\times\mathcal{X}\rightarrow$. The utility functions are defined as follows:
@@ 5. PrivAgg Mechanism
(PrivAgg Mechanism - Global Sensitivity). The global PrivAgg mechanism $PrivAgg_{global}(x,\epsilon,u_{1},\ldots,u_{m},\mathcal{R})$ takes as input a database $x\in\mathcal{X}$, the privacy budget $\epsilon$, the utility functions $u_{1},u_{2},\ldots,u_{m}:\mathcal{R}\times\mathcal{X}\rightarrow$, the set of candidate solutions $\mathcal{R}$, and outputs a candidate $r\in\mathcal{R}$. The mechanism is defined as follows:
@@
(PrivAgg Mechanism - Local Sensitivity). The local PrivAgg mechanism $PrivAgg_{local}(x,\epsilon,u_{1},\ldots,u_{m},\delta^{u_{1}},\ldots,\delta^{u_{m}},\mathcal{R})$ takes as input a database $x\in\mathcal{X}$, the privacy budget $\epsilon$, the set of utility functions $u_{1},u_{2},\ldots,u_{m}:\mathcal{R}\times\mathcal{X}\rightarrow$, the admissible sensitivity functions $\delta^{u_{1}},\ldots,\delta^{u_{m}}$ of the utility functions, a set of candidate solutions $\mathcal{R}$, and outputs a candidate $r\in\mathcal{R}$. The mechanism is defined as follows:
@@ 5.1. Sensitivity Analysis
. Let $u_{1},\ldots,u_{m}:\mathcal{R}\times\mathcal{X}\rightarrow$ be utility functions and $w_{1},\ldots,w_{m}$ be weights. We have
@@
Let $u_{1},\ldots,u_{m}:\mathcal{R}\times\mathcal{X}\rightarrow$ be utility functions and $w_{1},\ldots,w_{m}$ be weights. We have that
@@
Let $u_{1},\ldots,u_{m}:\mathcal{R}\times\mathcal{X}\rightarrow$ be utility functions, $\delta^{u_{1}},\ldots,\delta^{u_{m}}$ be admissible sensitivity functions, and $w_{1},\ldots,w_{m}$ be weights. Then an admissible sensitivity of the aggregate utility function $u_{agg}(x,r)$ is given by
@@ 6.2.3. Sensitivity Analysis
$P(x)$ is the number os positive samples in the database $x$ and $TP(x,\tau)$ is the number of true positive predictions of the tree $\tau$ on the database $x$. The sensitivity function for $\delta^{TNR}$ is defined similarly by replacing $TP$ by $TN$ and $P$ by $N$ where $TN(x,\tau)$ is the number of true negative predictions of the tree $\tau$ on the database $x$ and $N(x)$ is the number of negative samples in the database $x$. We show the following Lemma:

### v2

@@ Abstract.
Differentially private selection mechanisms are fundamental building blocks for privacy-preserving data analysis. While numerous mechanisms exist for single-objective selection, many real-world applications require optimizing multiple competing objectives simultaneously. We present two novel mechanisms for differentially private multi-objective selection: PrivPareto and PrivAgg. PrivPareto uses a novel Pareto score to identify solutions near the Pareto frontier, while PrivAgg enables privacy-preserving weighted aggregation of multiple objectives. Both mechanisms support global and local sensitivity approaches, with comprehensive theoretical analysis showing how to compose sensitivities of multiple utility functions. We demonstrate their practical applicability through two real-world applications: cost-sensitive decision tree construction and multi-objective influential node selection in social networks. The experimental results showed that our local sensitivity-based approaches achieve significantly better utility compared to global sensitivity approaches across both applications and both Pareto and Aggregation approaches. Moreover, the local sensitivity-based approaches are able to perform well with typical privacy budget values $\epsilon\in[0.01,1]$ in most experiments.
@@ 1. Introduction
Data analysis is a fundamental tool for decision making in many areas, such as healthcare, finance, and marketing. Data analysts can use data to extract valuable insights, make predictions, optimize decisions and publish their findings. On the other hand, a malicious adversary can use the published data to infer sensitive information about individuals (Henriksen-Bulmer and Jeary, 2016; Hamza et al., 2013; Narayanan and Shmatikov, 2016).
@@
Differential privacy has been widely applied to the problem of single-objective private selection. In this problem, the goal is to select an element from a set of candidates that maximizes a single objective function. Formally, given a set of candidates $\mathcal{R}$ and a utility function $u:\mathcal{X}\times\mathcal{R}\rightarrow\real$ that takes a candidate $r\in\mathcal{R}$ and a dataset $x\in\mathcal{X}$ and outputs a score, the goal is to select an element $r\in\mathcal{R}$ that approximately maximizes $u(x,r)$ while satisfying differential privacy. The utility function $u$ is designed by the data analyst and measures how well a candidate $r$ satisfies the objective of a given task. For the rest of the paper, we will use utility function and objective function interchangeably.
Many algorithms have been proposed to solve this problem under differential privacy: the exponential mechanism (McSherry and Talwar, 2007), the report noisy max algorithm (Dwork et al., 2014), the permute-and-flip mechanism (McKenna and Sheldon, 2020) and the local dampening mechanism (de Farias et al., 2020; Farias et al., 2023). These mechanisms use the notion of sensitivity in two flavors, global sensitivity and local sensitivity, where these notions measure how much the utility function changes when a single individual is added or removed from a dataset.
Such algorithms are very important in the field of differential privacy since they are fundamental building blocks and can be composed to create more complex algorithms. Their applications include private release of streams, histograms, databases and graphs (Zhu et al., 2017), data mining (Fletcher and Islam, 2019), machine learning (Ji et al., 2014; Liu et al., 2021) and social network analysis (Task and Clifton, 2012; Jiang et al., 2021).
@@
Thus, we define our problem statement of multi-objective private selection as follows. Given a set of candidates $\mathcal{R}$ and a set of utility functions $u_{1},u_{2},\ldots,u_{m}:\mathcal{X}\times\mathcal{R}\rightarrow\real$, the goal is to select an element $r\in\mathcal{R}$ that maximizes $\{u_{1}(x,r),u_{2}(x,r),\ldots,u_{m}(x,r)\}$ while satisfying differential privacy.
Note that the utility functions may be conflicting and there could be no candidate that maximizes all utility functions simultaneously. Instead, it is desired to find a set of candidates that are not dominated in all utility functions. These are said to be Pareto optimal and the set of all Pareto optimal candidates is called the Pareto front. In the example of decision trees, the Pareto front would contains trees such that no other tree is better in both true positive rate and true negative rate.
@@
To show the effectiveness and generality of our mechanisms, we conduct an extensive experimental evaluation on two distinct problems: (i) multi-objective influential node selection in social networks where we select the most influencial nodes in a graph based on multiple centrality measures, and (ii) cost-sensitive decision tree construction where we build a set of approximate Pareto optimal decision trees maximizing true positive rate and true negative rate under differential privacy.
@@ 2.1. Basic Definitions
($\epsilon$-Differential Privacy (Dwork et al., 2006a, b)). A randomized algorithm $\mathcal{M}$ satisfies $\epsilon$-differential privacy if, for any two databases $x$ and $y$ such that $d(x,y)\leq 1$, and for any possible output $O$ of $\mathcal{M}$, the following inequality holds:
@@ 2.3. Sensitivity Functions
Computing local sensitivity $LS^{u}(x,t)$ or element local sensitivity $LS^{u}(x,t,r)$ can be infeasible is some cases where it is an NP-hard problem (Nissim et al., 2007; Zhang et al., 2015). To overcome this, we can build a computationally efficient function $\delta^{u}(x,t,r)$ that computes as an upper bound for $LS^{u}(x,t)$ or $LS^{u}(x,t,r)$, while staying below $\Delta u$. These functions are denoted as sensitivity functions. A sensitivity function needs to have the following signature: $\delta^{u}:\mathcal{X}\times\mathbb{N}\times\mathcal{R}\rightarrow\real$ and the function $\delta^{u}$ is said to be a sensitivity function for the utility function $u$. The $\delta^{u}(x,t,r)=\Delta u$, $\delta^{u}(x,t,r)=LS^{u}(x,t)$ and $\delta^{u}(x,t,r)=LS^{u}(x,t,r)$ are all sensitivity functions.
@@ 3. Related Work
The field of differentially private multi-objective selection remains unexplored. To the best of our knowledge, this work is the first to tackle this problem. Nonetheless, we review the literature on related works in differential privacy for single-objective selection.
@@ 4. PrivPareto Mechanism
The choice of the single objective differentially private selection mechanism in the third step plays an important role on the accuracy of the our mechanism since each mechanism adds a different amount of noise to the output. These mechanisms can use global or local sensitivity. Global sensitivity based algorithms are more frequent in the literature (McSherry and Talwar, 2007; McKenna and Sheldon, 2020). However, local sensitivity based mechanisms have been shown to perform better in applications where the local sensitivity is smaller than the global sensitivity (Farias et al., 2023) which is the case for the Pareto Score. Thus, we present the PrivPareto mechanism in two versions: a global sensitivity version and a local sensitivity version.
@@
(Global PrivPareto Mechanism) The global PrivPareto mechanism $PrivPareto_{global}(x,\epsilon,u_{1},\ldots,u_{m},\mathcal{R})$ takes as input a database $x\in\mathcal{X}$, the privacy budget $\epsilon$, the utility functions $u_{1},u_{2},\ldots,u_{m}:\mathcal{X}\times\mathcal{R}\rightarrow\real$, the set of candidate solutions $\mathcal{R}$, and outputs an approximate Pareto optimal solution $r\in\mathcal{R}$. The mechanism is defined as follows:
@@
The local sensitivity version of the PrivPareto mechanism is given in Definition 4.2:
@@
(Local PrivPareto Mechanism) The local PrivPareto mechanism $PrivPareto_{local}(x,\epsilon,u_{1},\ldots,u_{m},\delta^{u_{1}},\ldots,\delta^{u_{m}},\mathcal{R})$ takes as input a database $x\in\mathcal{X}$, the privacy budget $\epsilon$, the set of utility functions $u_{1},u_{2},\ldots,u_{m}:\mathcal{X}\times\mathcal{R}\rightarrow\real$, the admissible sensitivity functions $\delta^{u_{1}},\ldots,\delta^{u_{m}}$ of the utility functions, a set of candidate solutions $\mathcal{R}$, and outputs an approximate Pareto optimal solution $r\in\mathcal{R}$. The mechanism is defined as follows:
@@ 4.1. Pareto Score
(Pareto Score). Given a database $x\in\mathcal{X}$, a set of utility functions $u_{1},u_{2},\ldots,u_{m}:\mathcal{X}\times\mathcal{R}\rightarrow\real$, and a candidate solution $r\in\mathcal{R}$, the Pareto Score $PS(x,r)$ is defined as the number of candidates that dominate $r$ in the set $\mathcal{R}$:
@@
Consider a database $x$, a set of candidates =$\mathcal{R}=\{a,b,c,d,e\}$ and two utility functions $u_{1},u_{2}:\mathcal{X}\times\mathcal{R}\rightarrow\real$. The utility functions are defined as follows:
@@
Figure 1. Example Pareto scores for 5 elements and 2 utility functions. The left lower subspace of each candidate, delimited by the two dashed lines, contains its dominated elements. For instance, candidate $a$ dominates $d$ and $e$.
@@ 4.2. Pareto Score Sensitivity
Consider a database $x$, a set of candidates =$\mathcal{R}=\{a,b,c\}$ and two utility functions $u_{1},u_{2}:\mathcal{X}\times\mathcal{R}\rightarrow\real$. The utility functions are defined as follows:
@@ 5. PrivAgg Mechanism
(PrivAgg Mechanism - Global Sensitivity). The global PrivAgg mechanism $PrivAgg_{global}(x,\epsilon,u_{1},\ldots,u_{m},\mathcal{R})$ takes as input a database $x\in\mathcal{X}$, the privacy budget $\epsilon$, the utility functions $u_{1},u_{2},\ldots,u_{m}:\mathcal{X}\times\mathcal{R}\rightarrow\real$, the set of candidate solutions $\mathcal{R}$, and outputs a candidate $r\in\mathcal{R}$. The mechanism is defined as follows:
@@
(PrivAgg Mechanism - Local Sensitivity). The local PrivAgg mechanism $PrivAgg_{local}(x,\epsilon,u_{1},\ldots,u_{m},\delta^{u_{1}},\ldots,\delta^{u_{m}},\mathcal{R})$ takes as input a database $x\in\mathcal{X}$, the privacy budget $\epsilon$, the set of utility functions $u_{1},u_{2},\ldots,u_{m}:\mathcal{X}\times\mathcal{R}\rightarrow\real$, the admissible sensitivity functions $\delta^{u_{1}},\ldots,\delta^{u_{m}}$ of the utility functions, a set of candidate solutions $\mathcal{R}$, and outputs a candidate $r\in\mathcal{R}$. The mechanism is defined as follows:
@@ 5.1. Sensitivity Analysis
. Let $u_{1},\ldots,u_{m}:\mathcal{X}\times\mathcal{R}\rightarrow\real$ be utility functions and $w_{1},\ldots,w_{m}$ be weights. We have
@@
Let $u_{1},\ldots,u_{m}:\mathcal{X}\times\mathcal{R}\rightarrow\real$ be utility functions and $w_{1},\ldots,w_{m}$ be weights. We have that
@@
Let $u_{1},\ldots,u_{m}:\mathcal{X}\times\mathcal{R}\rightarrow\real$ be utility functions, $\delta^{u_{1}},\ldots,\delta^{u_{m}}$ be admissible sensitivity functions, and $w_{1},\ldots,w_{m}$ be weights. Then an admissible sensitivity of the aggregate utility function $u_{agg}(x,r)$ is given by
@@ 6.2.3. Sensitivity Analysis
$P(x)$ is the number of positive samples in the database $x$ and $TP(x,\tau)$ is the number of true positive predictions of the tree $\tau$ on the database $x$. The sensitivity function for $\delta^{TNR}$ is defined similarly by replacing $TP$ by $TN$ and $P$ by $N$ where $TN(x,\tau)$ is the number of true negative predictions of the tree $\tau$ on the database $x$ and $N(x)$ is the number of negative samples in the database $x$. We show the following Lemma: